A think stick 1.5m long is pivoted about one end. The stick is held in a horizontal position and released from rest. Neglecting friction, use conservation of energy to determine the stick's angular velocity when it is vertical. (Hint: When determining the change in gravitational potential energy of the stick, you must consider the change in position of the stick's centre of mass)
Answer #1

Since stick is at rest initially , therefore
Initial energy = 0
Now when stick becomes vertical then
Final energy = potential energy + Kinetic energy
= mg(-L/2) + (1/2)Iw2
where I is moment of inertia of stick = mL2/3
L is length of stick
m is mass of stick
w is angular velocity
Now taking energy conservation
mg(-L/2) + (1/2)Iw2 = 0
(1/2)Iw2 = mg(L/2)
(1/2)(mL2/3)w2 = mg(L/2)
2
On putting the value we get
3*9.81 1.5 u2
w = 4.429 rad/s
LT